magnitude and direction of the magnetic fieldImage for 2. A 0.3-m long wire of 8.7g carries 10 A of current as shown. What is the magnitude and direction of the fiel
Answer #1

given,

mass of wire = 8.7 gm or 8.7 * 10^-3 kg

current = 10 A

length of wire = 0.3 m

to levitate the wire force due to magnetic force should be equal to force due to gravity

force due to magnetic force = magnetic field * length * current

magnetic field * length * current = mass * g

magnetic field * 0.3 * 10 = 8.7 * 10^-3 * 9.8

magnetic field = 0.02842 T

direction will be into the page